### v2

# On the semigroup of monoid endomorphisms of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ with a two-element family $\mathscr{F}$ of inductive nonempty subsets of $\omega$

Abstract.
We study the semigroup of non-injective monoid endomorphisms of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ with a two-elements family $\mathscr{F}$ of inductive nonempty subsets of $\omega$. We describe the structure of elements of the semigroup $\boldsymbol{End}^{*}_{0}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ of non-injective monoid endomorphisms of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$. In particular we show that its subsemigroup $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ of non-injective non-annihilating monoid endomorphisms of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ is isomorphic to the direct product of the two-element left-zero semigroup and the multiplicative semigroup of positive integers and describe Green’s relations on $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$.
Key words and phrases:
Bicyclic monoid, inverse semigroup, bicyclic extension, monoid endomorphism, non-injective, Green’s relations, left-zero semigroup, direct product
2020 Mathematics Subject Classification:
20M10, 20M12, 20M15, 20M18. 20M20.
We shall follow the terminology of [1, 2, 9]. By $\omega$ we denote the set of all non-negative integers, by $\mathbb{N}$ the set of all positive integers, and by $\mathbb{Z}$ the set of all integers.
Let $\mathscr{P}(\omega)$ be the family of all subsets of $\omega$. For any $F\in\mathscr{P}(\omega)$ and $n\in\mathbb{Z}$ we put $n+F=\{n+k\colon k\in F\}$ if $F\neq\varnothing$ and $n+\varnothing=\varnothing$. A subfamily $\mathscr{F}\subseteq\mathscr{P}(\omega)$ is called ${\omega}$-closed if $F_{1}\cap(-n+F_{2})\in\mathscr{F}$ for all $n\in\omega$ and $F_{1},F_{2}\in\mathscr{F}$. For any $a\in\omega$ we denote $[a)=\{x\in\omega\colon x\geqslant a\}$.
A subset $A$ of $\omega$ is said to be inductive, if $i\in A$ implies $i+1\in A$. Obviously, $\varnothing$ is an inductive subset of $\omega$.
Remark 1 ([5]).
  1. (1)
    By Lemma 6 from [4] a nonempty subset $F\subseteq\omega$ is inductive in $\omega$ if and only $(-1+F)\cap F=F$.
  2. (2)
    Since the set $\omega$ with the usual order is well-ordered, for any nonempty inductive subset $F$ in $\omega$ there exists a nonnegative integer $n_{F}\in\omega$ such that $[n_{F})=F$.
  3. (3)
    Statement (2) implies that the intersection of an arbitrary finite family of nonempty inductive subsets in $\omega$ is a nonempty inductive subset of $\omega$.
A semigroup $S$ is called inverse if for any element $x\in S$ there exists a unique $x^{-1}\in S$ such that $xx^{-1}x=x$ and $x^{-1}xx^{-1}=x^{-1}$. The element $x^{-1}$ is called the inverse of $x\in S$. If $S$ is an inverse semigroup, then the function $\operatorname{inv}\colon S\to S$ which assigns to every element $x$ of $S$ its inverse element $x^{-1}$ is called the inversion.
[…]
If $S$ is a semigroup, then we shall denote the Green relations on $S$ by $\mathscr{R}$, $\mathscr{L}$, $\mathscr{J}$, $\mathscr{D}$ and $\mathscr{H}$ (see [1, Section 2.1]):
$\qquad a\mathscr{R}b\mbox{ if and only if }aS^{1}=bS^{1};$
$\qquad a\mathscr{L}b\mbox{ if and only if }S^{1}a=S^{1}b;$
$\qquad a\mathscr{J}b\mbox{ if and only if }S^{1}aS^{1}=S^{1}bS^{1};$
$\qquad\mathscr{D}=\mathscr{L}\circ\mathscr{R}=\mathscr{R}\circ\mathscr{L};$
$\qquad\mathscr{H}=\mathscr{L}\cap\mathscr{R}.$
The $\mathscr{L}$-class [$\mathscr{R}$-class, $\mathscr{H}$-class, $\mathscr{D}$-class, $\mathscr{J}$-class] of the semigroup $S$ containing the element $a\in S$ will be denoted by $\boldsymbol{L}_{a}$ [$\boldsymbol{R}_{a}$, $\boldsymbol{H}_{a}$, $\boldsymbol{D}_{a}$, $\boldsymbol{J}_{a}$].
The bicyclic monoid ${\mathscr{C}}(p,q)$ is the semigroup with the identity $1$ generated by two elements $p$ and $q$ subjected only to the condition $pq=1$. The semigroup operation on ${\mathscr{C}}(p,q)$ is determined as follows:
$q^{k}p^{l}\cdot q^{m}p^{n}=q^{k+m-\min\{l,m\}}p^{l+n-\min\{l,m\}}.$
It is well known that the bicyclic monoid ${\mathscr{C}}(p,q)$ is a bisimple (and hence simple) combinatorial $E$-unitary inverse semigroup and every non-trivial congruence on ${\mathscr{C}}(p,q)$ is a group congruence [1].
On the set $\boldsymbol{B}_{\omega}=\omega\times\omega$ we define the semigroup operation $\cdot$” in the following way
(1) $(i_{1},j_{1})\cdot(i_{2},j_{2})=\left\{\begin{array}[]{ll}(i_{1}-j_{1}+i_{2},j_{2}),&\hbox{if~{}}j_{1}\leqslant i_{2};\\ (i_{1},j_{1}-i_{2}+j_{2}),&\hbox{if~{}}j_{1}\geqslant i_{2}.\end{array}\right.$
It is well known that the bicyclic monoid $\mathscr{C}(p,q)$ is isomorphic to the semigroup $\boldsymbol{B}_{\omega}$ by the mapping $\mathfrak{h}\colon\mathscr{C}(p,q)\to\boldsymbol{B}_{\omega}$, $q^{k}p^{l}\mapsto(k,l)$, $k,l\in\omega$ (see: [1, Section 1.12] or [11, Exercise IV.1.11$(ii)$]). Later we identify the bicyclic monoid $\mathscr{C}(p,q)$ with the semigroup $\boldsymbol{B}_{\omega}$ by the mapping $\mathfrak{h}$.
Next we shall describe the construction which is introduced in [4].
Let $\boldsymbol{B}_{\omega}$ be the bicyclic monoid and $\mathscr{F}$ be an ${\omega}$-closed subfamily of $\mathscr{P}(\omega)$. On the set $\boldsymbol{B}_{\omega}\times\mathscr{F}$ we define the semigroup operation $\cdot$” in the following way
(2) $(i_{1},j_{1},F_{1})\cdot(i_{2},j_{2},F_{2})=\left\{\begin{array}[]{ll}(i_{1}-j_{1}+i_{2},j_{2},(j_{1}-i_{2}+F_{1})\cap F_{2}),&\hbox{if~{}}j_{1}\leqslant i_{2};\\ (i_{1},j_{1}-i_{2}+j_{2},F_{1}\cap(i_{2}-j_{1}+F_{2})),&\hbox{if~{}}j_{1}\geqslant i_{2}.\end{array}\right.$
In [4] is proved that if the family $\mathscr{F}\subseteq\mathscr{P}(\omega)$ is ${\omega}$-closed then $(\boldsymbol{B}_{\omega}\times\mathscr{F},\cdot)$ is a semigroup. Moreover, if an ${\omega}$-closed family $\mathscr{F}\subseteq\mathscr{P}(\omega)$ contains the empty set $\varnothing$ then the set $\boldsymbol{I}=\{(i,j,\varnothing)\colon i,j\in\omega\}$ is an ideal of the semigroup $(\boldsymbol{B}_{\omega}\times\mathscr{F},\cdot)$. For any ${\omega}$-closed family $\mathscr{F}\subseteq\mathscr{P}(\omega)$ the following semigroup
$\boldsymbol{B}_{\omega}^{\mathscr{F}}=\left\{\begin{array}[]{ll}(\boldsymbol{B}_{\omega}\times\mathscr{F},\cdot)/\boldsymbol{I},&\hbox{if~{}}\varnothing\in\mathscr{F};\\ (\boldsymbol{B}_{\omega}\times\mathscr{F},\cdot),&\hbox{if~{}}\varnothing\notin\mathscr{F}\end{array}\right.$
is defined in [4]. The semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ generalizes the bicyclic monoid and the countable semigroup of matrix units. It is proven in [4] that $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ is a combinatorial inverse semigroup and Green’s relations, the natural partial order on $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ and its set of idempotents are described. Here, the criteria when the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ is simple, 00-simple, bisimple, 00-bisimple, or it has the identity, are given. In particularly in [4] it is proved that the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ is isomorphic to the semigrpoup of ${\omega}{\times}{\omega}$-matrix units if and only if $\mathscr{F}$ consists of a singleton set and the empty set, and $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ is isomorphic to the bicyclic monoid if and only if $\mathscr{F}$ consists of a non-empty inductive subset of $\omega$.
Group congruences on the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ and its homomorphic retracts in the case when an ${\omega}$-closed family $\mathscr{F}$ consists of inductive non-empty subsets of $\omega$ are studied in [5]. It is proven that a congruence $\mathfrak{C}$ on $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ is a group congruence if and only if its restriction on a subsemigroup of $\boldsymbol{B}_{\omega}^{\mathscr{F}}$, which is isomorphic to the bicyclic semigroup, is not the identity relation. Also in [5], all non-trivial homomorphic retracts and isomorphisms of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ are described. In [6] it is proved that an injective endomorphism $\varepsilon$ of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ is the indentity transformation if and only if $\varepsilon$ has three distinct fixed points, which is equivalent to existence non-idempotent element $(i,j,[p))\in\boldsymbol{B}_{\omega}^{\mathscr{F}}$ such that $(i,j,[p))\varepsilon=(i,j,[p))$.
In [3, 10] the algebraic structure of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ is established in the case when ${\omega}$-closed family $\mathscr{F}$ consists of atomic subsets of ${\omega}$.
It is well-known that every automorphism of the bicyclic monoid $\boldsymbol{B}_{\omega}$ is the identity self-map of $\boldsymbol{B}_{\omega}$ [1], and hence the group $\mathbf{Aut}(\boldsymbol{B}_{\omega})$ of automorphisms of $\boldsymbol{B}_{\omega}$ is trivial. In [8] it is proved that the semigroup $\mathrm{\mathbf{End}}(\boldsymbol{B}_{\omega})$ of all endomorphisms of the bicyclic semigroup $\boldsymbol{B}_{\omega}$ is isomorphic to the semidirect products $(\omega,+)\rtimes_{\varphi}(\omega,*)$, where $+$ and * are the usual addition and the usual multiplication on $\omega$.
In the paper [7] we study injective endomorphisms of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ with the two-elements family $\mathscr{F}$ of inductive nonempty subsets of $\omega$. We describe the elements of the semigroup $\boldsymbol{End}^{1}_{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ of all injective monoid endomorphisms of the monoid $\boldsymbol{B}_{\omega}^{\mathscr{F}}$. In particular we show that every element of the semigroup $\boldsymbol{End}^{1}_{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ has a form either $\alpha_{k,p}$ or $\beta_{k,p}$, where the endomorphism $\alpha_{k,p}$ is defined by the formulae
$(i,j,[0))\alpha_{k,p}$ $=(ki,kj,[0)),$
$(i,j,[1))\alpha_{k,p}$ $=(p+ki,p+kj,[1)),$
for an arbitrary positive integer $k$ and any $p\in\{0,\ldots,k-1\}$, and the endomorphism $\beta_{k,p}$ is defined by the formulae
$(i,j,[0))\beta_{k,p}$ $=(ki,kj,[0)),$
$(i,j,[1))\beta_{k,p}$ $=(p+ki,p+kj,[0)),$
an arbitrary positive integer $k\geqslant 2$ and any $p\in\{1,\ldots,k-1\}$. In [7] we describe the product of elements of the semigroup $\boldsymbol{End}^{1}_{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$:
$\alpha_{k_{1},p_{1}}\alpha_{k_{2},p_{2}}$ $=\alpha_{k_{1}k_{2},p_{2}+k_{2}p_{1}};$
$\alpha_{k_{1},p_{1}}\beta_{k_{2},p_{2}}$ $=\beta_{k_{1}k_{2},p_{2}+k_{2}p_{1}};$
$\beta_{k_{1},p_{1}}\beta_{k_{2},p_{2}}$ $=\beta_{k_{1}k_{2},k_{2}p_{1}};$
$\beta_{k_{1},p_{1}}\alpha_{k_{2},p_{2}}$ $=\beta_{k_{1}k_{2},k_{2}p_{1}}.$
Also, here we prove that Green’s relations $\mathscr{R}$, $\mathscr{L}$, $\mathscr{H}$, $\mathscr{D}$, and $\mathscr{J}$ on $\boldsymbol{End}^{1}_{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ coincide with the equality relation.
Later we assume that an ${\omega}$-closed family $\mathscr{F}$ consists of two nonempty inductive nonempty subsets of $\omega$.
This paper is a continuation of [7]. We study non-injective monoid endomorphisms of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$. We describe the structure of elements of the semigroup $\boldsymbol{End}_{0}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ of all non-injective monoid endomorphisms of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$. In particular we show that its subsemigroup $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ of all non-injective non-annihilating monoid endomorphisms of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ is isomorphic to the direct product the two-element left-zero semigroup and the multiplicative semigroup of positive integers and describe Green’s relations on $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$.
Remark 2.
By Proposition 1 of [5] for any $\omega$-closed family $\mathscr{F}$ of inductive subsets in $\mathscr{P}(\omega)$ there exists an $\omega$-closed family $\mathscr{F}^{*}$ of inductive subsets in $\mathscr{P}(\omega)$ such that $[0)\in\mathscr{F}^{*}$ and the semigroups $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ and $\boldsymbol{B}_{\omega}^{\mathscr{F}^{*}}$ are isomorphic. Hence without loss of generality we may assume that the family $\mathscr{F}$ contains the set $[0)$.
If $\mathscr{F}$ is an arbitrary $\omega$-closed family $\mathscr{F}$ of inductive subsets in $\mathscr{P}(\omega)$ and $[s)\in\mathscr{F}$ for some $s\in\omega$ then
$\boldsymbol{B}_{\omega}^{\{[s)\}}=\{(i,j,[s))\colon i,j\in\omega\}$
is a subsemigroup of $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ [5] and by Proposition 3 of [4] the semigroup $\boldsymbol{B}_{\omega}^{\{[s)\}}$ is isomorphic to the bicyclic semigroup.
Lemma 1.
Let =$\mathscr{F}=\{[0),[1)\}$ and let $\mathfrak{e}$ be a monoid endomorphism of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$. If $(i_{1},j_{1},F)\mathfrak{e}=(i_{2},j_{2},F)\mathfrak{e}$ for distinct two elements $(i_{1},j_{1},F),(i_{2},j_{2},F)$ of $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ for some $F\in\mathscr{F}$ then $\mathfrak{e}$ is the annihilating endomorphism of $\boldsymbol{B}_{\omega}^{\mathscr{F}}$.
Proof.
By Theorem 1 of [5] the image $(\boldsymbol{B}_{\omega}^{\mathscr{F}})\mathfrak{e}$ is a subgroup of $\boldsymbol{B}_{\omega}^{\mathscr{F}}$. By Theorem 4$(iii)$ of [4] every $\mathscr{H}$-class in $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ is a singleton, and hence $\mathfrak{e}$ is the annihilating monoid endomorphism of $\boldsymbol{B}_{\omega}^{\mathscr{F}}$. ∎
Lemma 2.
Let =$\mathscr{F}=\{[0),[1)\}$. Then $(\boldsymbol{B}_{\omega}^{\mathscr{F}})\mathfrak{e}\subseteq\boldsymbol{B}_{\omega}^{\{[0)\}}$ for any non-injective monoid endomorphism $\mathfrak{e}$ of $\boldsymbol{B}_{\omega}^{\mathscr{F}}$.
Proof.
By Proposition 3 of [4] the subsemigroup $\boldsymbol{B}_{\omega}^{\{[0)\}}$ of $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ is isomorphic to the bicyclic semigroup and hence by Corollary 1.32 of [1] the image $(\boldsymbol{B}_{\omega}^{\{[0)\}})\mathfrak{e}$ either is isomorphic to the bicyclic semigroup or is a cyclic subgroup of $\boldsymbol{B}_{\omega}^{\mathscr{F}}$. Since $(0,0,[0))\mathfrak{e}=(0,0,[0))$, Proposition 4 from [5] implies that $(\boldsymbol{B}_{\omega}^{\{[0)\}})\mathfrak{e}\subseteq\boldsymbol{B}_{\omega}^{\{[0)\}}$ in the case when the image $(\boldsymbol{B}_{\omega}^{\{[0)\}})\mathfrak{e}$ is isomorphic to the bicyclic semigroup. In the other case we have that the equality $(0,0,[0))\mathfrak{e}=(0,0,[0))$ implies that
$(\boldsymbol{B}_{\omega}^{\{[0)\}})\mathfrak{e}\subseteq\{(0,0,[0))\}\subseteq\boldsymbol{B}_{\omega}^{\{[0)\}},$
because by Theorem 4$(iii)$ of [4] every $\mathscr{H}$-class in $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ is a singleton.
Next, by Proposition 3 of [4] the subsemigroup $\boldsymbol{B}_{\omega}^{\{[1)\}}$ of $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ is isomorphic to the bicyclic semigroup and hence by Corollary 1.32 of [1] the image $(\boldsymbol{B}_{\omega}^{\{[1)\}})\mathfrak{e}$ either is isomorphic to the bicyclic semigroup or is a cyclic subgroup of $\boldsymbol{B}_{\omega}^{\mathscr{F}}$. Suppose that the image $(\boldsymbol{B}_{\omega}^{\{[1)\}})\mathfrak{e}$ is isomorphic to the bicyclic semigroup and $(\boldsymbol{B}_{\omega}^{\{[1)\}})\mathfrak{e}\subseteq\boldsymbol{B}_{\omega}^{\{[1)\}}$. Then monoid endomorphism $\mathfrak{e}$ of $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ is injective. Indeed, injectivity of the restriction $\mathfrak{e}{\upharpoonleft}_{\boldsymbol{B}_{\omega}^{\{[1)\}}}\boldsymbol{B}_{\omega}^{\{[1)\}}\to\boldsymbol{B}_{\omega}^{\{[1)\}}$, Proposition 4 of [5], Corollary 1.32 of [1], Theorem 4$(iii)$ of [4], and the equality $(0,0,[0))\mathfrak{e}=(0,0,[0))$ imply that either the restriction $\mathfrak{e}{\upharpoonleft}_{\boldsymbol{B}_{\omega}^{\{[0)\}}}\boldsymbol{B}_{\omega}^{\{[0)\}}\to\boldsymbol{B}_{\omega}^{\{[0)\}}$ is an injective mapping or is an annihilating endomorphism. In the case when the restriction $\mathfrak{e}{\upharpoonleft}_{\boldsymbol{B}_{\omega}^{\{[0)\}}}\boldsymbol{B}_{\omega}^{\{[0)\}}\to\boldsymbol{B}_{\omega}^{\{[0)\}}$ is an injective mapping we get that the endomorphism $\mathfrak{e}$ is injective. If the image $(\boldsymbol{B}_{\omega}^{\{[0)\}})\mathfrak{e}$ is a singleton then by Lemma 1 we have that $\mathfrak{e}$ is the annihilating monoid endomorphism of $\boldsymbol{B}_{\omega}^{\mathscr{F}}$. In the both cases we obtain that $(\boldsymbol{B}_{\omega}^{\mathscr{F}})\mathfrak{e}\subseteq\boldsymbol{B}_{\omega}^{\{[0)\}}$. ∎
Example 1.
Let =$\mathscr{F}=\{[0),[1)\}$ and $k$ be an arbitrary non-negative integer. We define a map $\gamma_{k}\colon\boldsymbol{B}_{\omega}^{\mathscr{F}}\to\boldsymbol{B}_{\omega}^{\mathscr{F}}$ by the formulae
$(i,j,[0))\gamma_{k}=(i,j,[1))\gamma_{k}=(ki,kj,[0))$
for all $i,j\in\omega$.
We claim that $\gamma_{k}\colon\boldsymbol{B}_{\omega}^{\mathscr{F}}\to\boldsymbol{B}_{\omega}^{\mathscr{F}}$ is an endomorphism. Example 2 and Proposition 5 from [5] imply that the map $\gamma_{1}\colon\boldsymbol{B}_{\omega}^{\mathscr{F}}\to\boldsymbol{B}_{\omega}^{\mathscr{F}}$ is a homomorphic retraction of the monoid $\boldsymbol{B}_{\omega}^{\mathscr{F}}$, and hence it is a monoid endomorphism of $\boldsymbol{B}_{\omega}^{\mathscr{F}}$. By Lemma 2 of [8] every monoid endomorphism $\mathfrak{h}$ of the semigroup $\boldsymbol{B}_{\omega}$ has the following form
$(i,j)\mathfrak{h}=(ki,kj),\qquad\hbox{for some}\quad k\in\omega.$
This implies that the map $\gamma_{k}$ is a monoid endomorphism of $\boldsymbol{B}_{\omega}^{\mathscr{F}}$.
Example 2.
Let =$\mathscr{F}=\{[0),[1)\}$ and $k$ be an arbitrary non-negative integer. We define a map $\delta_{k}\colon\boldsymbol{B}_{\omega}^{\mathscr{F}}\to\boldsymbol{B}_{\omega}^{\mathscr{F}}$ by the formulae
$(i,j,[0))\delta_{k}=(ki,kj,[0))\qquad\hbox{and}\qquad(i,j,[1))\delta_{k}=(k(i+1),k(j+1),[0))$
for all $i,j\in\omega$.
Proposition 1.
Let =$\mathscr{F}=\{[0),[1)\}$. Then for any $k\in\omega$ the map $\delta_{k}$ is an endomorphism of the monoid $\boldsymbol{B}_{\omega}^{\mathscr{F}}$.
Proof.
Since by Proposition 3 of [4] the subsemigroups $\boldsymbol{B}_{\omega}^{\{[0)\}}$ and $\boldsymbol{B}_{\omega}^{\{[1)\}}$ of $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ are isomorphic to the bicyclic semigroup, by Lemma 2 of [8] the restrictions $\delta_{k}{\upharpoonleft}_{\boldsymbol{B}_{\omega}^{\{[0)\}}}\colon\boldsymbol{B}_{\omega}^{\{[0)\}}$ $\to\boldsymbol{B}_{\omega}^{\mathscr{F}}$ and $\delta_{k}{\upharpoonleft}_{\boldsymbol{B}_{\omega}^{[1)}}\colon\boldsymbol{B}_{\omega}^{\{[1)\}}\to\boldsymbol{B}_{\omega}^{\mathscr{F}}$ of $\delta_{k}$ are homomorphisms. Hence it sufficient to show that the following equalities
$(i_{1},j_{1},[0))\delta_{k}\cdot(i_{2},j_{2},[1))\delta_{k}$ $=((i_{1},j_{1},[0))\cdot(i_{2},j_{2},[1)))\delta_{k};$
$(i_{1},j_{1},[1))\delta_{k}\cdot(i_{2},j_{2},[0))\delta_{k}$ $=((i_{1},j_{1},[1))\cdot(i_{2},j_{2},[0)))\delta_{k},$
hold for any $i_{1},j_{1},i_{2},j_{2}\in\omega$.
We observe that the above equalities are trivial in the case when $k=0$. Hence later we assume that $k$ is a positive integer.
Then we have that
$(i_{1},$ $j_{1},[0))\delta_{k}\cdot(i_{2},j_{2},[1))\delta_{k}=(ki_{1},kj_{1},[0))\cdot(k(i_{2}+1),k(j_{2}+1),[0))=$
=$=\left\{\begin{array}[]{ll}(ki_{1}{+}k(i_{2}{+}1){-}kj_{1},k(j_{2}{+}1),(kj_{1}{-}k(i_{2}{+}1){+}[0))\cap[0)),&\hbox{if~{}}kj_{1}<k(i_{2}+1);\\ (ki_{1},k(j_{2}+1),[0)\cap[0)),&\hbox{if~{}}kj_{1}=k(i_{2}+1);\\ (ki_{1},kj_{1}{+}k(j_{2}{+}1){-}k(i_{2}{+}1),[0)\cap(k(i_{2}{+}1){-}kj_{1}{+}[0))),&\hbox{if~{}}kj_{1}>k(i_{2}+1)\end{array}\right.$
=$=\left\{\begin{array}[]{ll}(k(i_{1}+i_{2}+1-j_{1}),k(j_{2}+1),[0)),&\hbox{if~{}}j_{1}<i_{2}+1;\\ (ki_{1},k(j_{2}+1),[0)),&\hbox{if~{}}j_{1}=i_{2}+1;\\ (ki_{1},k(j_{1}+j_{2}-i_{2}),[0)),&\hbox{if~{}}j_{1}>i_{2}+1\end{array}\right.$
=$=\left\{\begin{array}[]{ll}(k(i_{1}+i_{2}+1-j_{1}),k(j_{2}+1),[0)),&\hbox{if~{}}j_{1}<i_{2};\\ (k(i_{1}+i_{2}+1-j_{1}),k(j_{2}+1),[0)),&\hbox{if~{}}j_{1}=i_{2};\\ (ki_{1},k(j_{2}+1),[0)),&\hbox{if~{}}j_{1}=i_{2}+1;\\ (ki_{1},k(j_{1}+j_{2}-i_{2}),[0)),&\hbox{if~{}}j_{1}>i_{2}+1\end{array}\right.$
=$=\left\{\begin{array}[]{ll}(k(i_{1}+i_{2}+1-j_{1}),k(j_{2}+1),[0)),&\hbox{if~{}}j_{1}<i_{2};\\ (k(i_{1}+1),k(j_{2}+1),[0)),&\hbox{if~{}}j_{1}=i_{2};\\ (ki_{1},k(j_{2}+1),[0)),&\hbox{if~{}}j_{1}=i_{2}+1;\\ (ki_{1},k(j_{1}+j_{2}-i_{2}),[0)),&\hbox{if~{}}j_{1}>i_{2}+1,\end{array}\right.$
$((i_{1},j_{1},[0))\cdot(i_{2},j_{2},[1)))\delta_{k}$ =$=\left\{\begin{array}[]{ll}(i_{1}+i_{2}-j_{1},j_{2},(j_{1}-i_{2}+[0))\cap[1))\delta_{k},&\hbox{if~{}}j_{1}<i_{2};\\ (i_{1},j_{2},[0)\cap[1))\delta_{k},&\hbox{if~{}}j_{1}=i_{2};\\ (i_{1},j_{1}+j_{2}-i_{2},[0)\cap(i_{2}-j_{1}+[1)))\delta_{k},&\hbox{if~{}}j_{1}>i_{2}\end{array}\right.$
=$=\left\{\begin{array}[]{ll}(i_{1}+i_{2}-j_{1},j_{2},[1))\delta_{k},&\hbox{if~{}}j_{1}<i_{2};\\ (i_{1},j_{2},[1))\delta_{k},&\hbox{if~{}}j_{1}=i_{2};\\ (i_{1},j_{1}+j_{2}-i_{2},[0))\delta_{k},&\hbox{if~{}}j_{1}>i_{2}\end{array}\right.$
=$=\left\{\begin{array}[]{ll}(k(i_{1}+i_{2}-j_{1}+1),k(j_{2}+1),[0)),&\hbox{if~{}}j_{1}<i_{2};\\ (k(i_{1}+1),k(j_{2}+1),[0)),&\hbox{if~{}}j_{1}=i_{2};\\ (ki_{1},k(j_{1}+j_{2}-i_{2}),[0)),&\hbox{if~{}}j_{1}>i_{2}\end{array}\right.$
=$=\left\{\begin{array}[]{ll}(k(i_{1}+i_{2}-j_{1}+1),k(j_{2}+1),[0)),&\hbox{if~{}}j_{1}<i_{2};\\ (k(i_{1}+1),k(j_{2}+1),[0)),&\hbox{if~{}}j_{1}=i_{2};\\ (ki_{1},k(j_{1}+j_{2}-i_{2}),[0)),&\hbox{if~{}}j_{1}=i_{2}+1;\\ (ki_{1},k(j_{1}+j_{2}-i_{2}),[0)),&\hbox{if~{}}j_{1}>i_{2}+1\end{array}\right.$
=$=\left\{\begin{array}[]{ll}(k(i_{1}+i_{2}+1-j_{1}),k(j_{2}+1),[0)),&\hbox{if~{}}j_{1}<i_{2};\\ (k(i_{1}+1),k(j_{2}+1),[0)),&\hbox{if~{}}j_{1}=i_{2};\\ (ki_{1},k(j_{2}+1),[0)),&\hbox{if~{}}j_{1}=i_{2}+1;\\ (ki_{1},k(j_{1}+j_{2}-i_{2}),[0)),&\hbox{if~{}}j_{1}>i_{2}+1,\end{array}\right.$
and
$(i_{1},j_{1},$ $[1))\delta_{k}\cdot(i_{2},j_{2},[0))\delta_{k}=(k(i_{1}+1),k(j_{1}+1),[0))\cdot(ki_{2},kj_{2},[0))=$
=$=\left\{\begin{array}[]{ll}(k(i_{1}{+}1){+}ki_{2}{-}k(j_{1}{+}1),kj_{2},(k(j_{1}{+}1){-}ki_{2}{+}[0))\cap[0)),&\hbox{if~{}}k(j_{1}+1)<ki_{2};\\ (k(i_{1}+1),kj_{2},[0)\cap[0)),&\hbox{if~{}}k(j_{1}+1)=ki_{2};\\ (k(i_{1}{+}1),k(j_{1}{+}1){+}kj_{2}{-}ki_{2},[0)\cap(ki_{2}{-}k(j_{1}{+}1){+}[0))),&\hbox{if~{}}k(j_{1}+1)>ki_{2}\end{array}\right.$
=$=\left\{\begin{array}[]{ll}(k(i_{1}+i_{2}-j_{1}),kj_{2},[0)),&\hbox{if~{}}j_{1}+1<i_{2};\\ (k(i_{1}+1),kj_{2},[0)),&\hbox{if~{}}j_{1}+1=i_{2};\\ (k(i_{1}+1),k(j_{1}+1+j_{2}-i_{2}),[0)),&\hbox{if~{}}j_{1}+1>i_{2}\end{array}\right.$
=$=\left\{\begin{array}[]{ll}(k(i_{1}+i_{2}-j_{1}),kj_{2},[0)),&\hbox{if~{}}j_{1}+1<i_{2};\\ (k(i_{1}+1),kj_{2},[0)),&\hbox{if~{}}j_{1}+1=i_{2};\\ (k(i_{1}+1),k(j_{1}+1+j_{2}-i_{2}),[0)),&\hbox{if~{}}j_{1}=i_{2};\\ (k(i_{1}+1),k(j_{1}+1+j_{2}-i_{2}),[0)),&\hbox{if~{}}j_{1}+1>i_{2}\end{array}\right.$
=$=\left\{\begin{array}[]{ll}(k(i_{1}+i_{2}-j_{1}),kj_{2},[0)),&\hbox{if~{}}j_{1}+1<i_{2};\\ (k(i_{1}+1),kj_{2},[0)),&\hbox{if~{}}j_{1}+1=i_{2};\\ (k(i_{1}+1),k(j_{2}+1),[0)),&\hbox{if~{}}j_{1}=i_{2};\\ (k(i_{1}+1),k(j_{1}+j_{2}-i_{2}+1),[0)),&\hbox{if~{}}j_{1}>i_{2},\end{array}\right.$
$((i_{1},j_{1},[1))\cdot(i_{2},j_{2},[0)))\delta_{k}$ =$=\left\{\begin{array}[]{ll}(i_{1}+i_{2}-j_{1},j_{2},(j_{1}-i_{2}+[1))\cap[0))\delta_{k},&\hbox{if~{}}j_{1}<i_{2};\\ (i_{1},j_{2},[1)\cap[0))\delta_{k},&\hbox{if~{}}j_{1}=i_{2};\\ (i_{1},j_{1}+j_{2}-i_{2},[1)\cap(i_{2}-j_{1}+[0)))\delta_{k},&\hbox{if~{}}j_{1}>i_{2}\end{array}\right.$
=$=\left\{\begin{array}[]{ll}(i_{1}+i_{2}-j_{1},j_{2},[0))\delta_{k},&\hbox{if~{}}j_{1}<i_{2};\\ (i_{1},j_{2},[1))\delta_{k},&\hbox{if~{}}j_{1}=i_{2};\\ (i_{1},j_{1}+j_{2}-i_{2},[1))\delta_{k},&\hbox{if~{}}j_{1}>i_{2}\end{array}\right.$
=$=\left\{\begin{array}[]{ll}(k(i_{1}+i_{2}-j_{1}),kj_{2},[0)),&\hbox{if~{}}j_{1}+1<i_{2};\\ (k(i_{1}+i_{2}-j_{1}),kj_{2},[0)),&\hbox{if~{}}j_{1}+1=i_{2};\\ (k(i_{1}+1),k(j_{2}+1),[0)),&\hbox{if~{}}j_{1}=i_{2};\\ (k(i_{1}+1),k(j_{1}+j_{2}-i_{2}+1),[0)),&\hbox{if~{}}j_{1}>i_{2}\end{array}\right.$
=$=\left\{\begin{array}[]{ll}(k(i_{1}+i_{2}-j_{1}),kj_{2},[0)),&\hbox{if~{}}j_{1}+1<i_{2};\\ (k(i_{1}+1),kj_{2},[0)),&\hbox{if~{}}j_{1}+1=i_{2};\\ (k(i_{1}+1),k(j_{2}+1),[0)),&\hbox{if~{}}j_{1}=i_{2};\\ (k(i_{1}+1),k(j_{1}+j_{2}-i_{2}+1),[0)),&\hbox{if~{}}j_{1}>i_{2}.\end{array}\right.$
This completes the proof of the statement of the proposition. ∎
Remark 3.
It obvious that if $\mathfrak{e}$ is the annihilating endomorphism of the monoid $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ then $\mathfrak{e}=\gamma_{0}=\delta_{0}$.
By $\boldsymbol{End}_{0}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ we denote the semigroup of all non-injective monoid endomorphisms of the monoid $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ for the family =$\mathscr{F}=\{[0),[1)\}$.
Theorems 1 and 2 describe the algebraic structure of the semigroup $\boldsymbol{End}_{0}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$.
Theorem 1.
If =$\mathscr{F}=\{[0),[1)\}$, then for any non-injective monoid endomorphism $\mathfrak{e}$ of the monoid $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ only one of the following conditions holds:
  1. (1)
    $\mathfrak{e}$ is the annihilating endomorphism, i.e., $\mathfrak{e}=\gamma_{0}=\delta_{0}$;
  2. (2)
    $\mathfrak{e}=\gamma_{k}$ for some positive integer $k$;
  3. (3)
    $\mathfrak{e}=\delta_{k}$ for some positive integer $k$.
Proof.
Fix an arbitrary non-injective monoid endomorphism $\mathfrak{e}$ of the monoid $\boldsymbol{B}_{\omega}^{\mathscr{F}}$. If $\mathfrak{e}$ is the annihilating endomorphism then statement (1) holds. Hence, later we assume that the endomorphism $\mathfrak{e}$ is not annihilating.
By Lemma 1 the restriction $\mathfrak{e}{\upharpoonleft}_{\boldsymbol{B}_{\omega}^{\{[0)\}}}\boldsymbol{B}_{\omega}^{\{[0)\}}\to\boldsymbol{B}_{\omega}^{\mathscr{F}}$ of the endomorphism $\mathfrak{e}$ is an injective mapping. Since by Proposition 3 of [4] the subsemigroup $\boldsymbol{B}_{\omega}^{\{[0)\}}$ of $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ are isomorphic to the bicyclic semigroup, the injectivity of the restriction $\mathfrak{e}{\upharpoonleft}_{\boldsymbol{B}_{\omega}^{\{[0)\}}}$ of the endomorphism $\mathfrak{e}$, Proposition 4 of [5], and Lemma 2 of [8] imply that there exists a positive integer $k$ such that
(3) $(i,j,[0))\mathfrak{e}=(ki,kj,[0)),$
for all $i,j\in\omega$.
By Lemma 1 the restriction $\mathfrak{e}{\upharpoonleft}_{\boldsymbol{B}_{\omega}^{\{[1)\}}}\boldsymbol{B}_{\omega}^{\{[1)\}}\to\boldsymbol{B}_{\omega}^{\mathscr{F}}$ of the endomorphism $\mathfrak{e}$ is an injective mapping, and by Lemma 2 we have that $(\boldsymbol{B}_{\omega}^{\{[1)\}})\mathfrak{e}\subseteq\boldsymbol{B}_{\omega}^{\{[0)\}}$. By Proposition 1.4.21$(6)$ of [9] a homomorphism of inverse semigroups preserves the natural partial order, and hence the following inequalities
[…]
Lemma 2, and Propositions 2 of [5] imply that
$(k,k,[0))$ $=(1,1,[0))\mathfrak{e}\preccurlyeq$
$\preccurlyeq(s,s,[0))=$
$=(0,0,[1))\mathfrak{e}\preccurlyeq$
[…]
$=(0,0,[0))\mathfrak{e}$
for some $s\in\{0,1,\ldots,k\}$. Again by Proposition 1.4.21$(6)$ of [9] and by Lemma 2 we get that
$(1,1,[1))\mathfrak{e}=(s+p,s+p,[0))$
for some non-negative integer $p$. If $p=0$ then $(1,1,[1))\mathfrak{e}=(0,0,[1))\mathfrak{e}$. By Lemma 1 the endomorphism $\mathfrak{e}$ is annihilating. Hence we assume that $p$ is a positive integer.
Let $(0,1,[1))\mathfrak{e}=(x,y,[0))$ for some $x,y\in\omega$. By Proposition $(1)$ of [9] and Lemma 4 of [4] we have that
$(1,0,[1))\mathfrak{e}$ $=((0,1,[1))^{-1})\mathfrak{e}=$
$=((0,1,[1))\mathfrak{e})^{-1}=$
$=(x,y,[0))^{-1}=$
$=(y,x,[0)).$
Since
[…]
the equalities $(0,0,[1))\mathfrak{e}=(s,s,[0))$ and $(1,1,[1))\mathfrak{e}=(s+p,s+p,[0))$ imply that
$(s,s,[0))$ $=(0,0,[1))\mathfrak{e}=$
$=((0,1,[1))\cdot(1,0,[1)))\mathfrak{e}=$
$=(0,1,[1))\mathfrak{e}\cdot(1,0,[1))\mathfrak{e}=$
$=(x,y,[0))\cdot(y,x,[0))=$
$=(x,x,[0))$
and
$(s+p,s+p,[0))$ $=(1,1,[1))\mathfrak{e}=$
$=((1,0,[1))\cdot(0,1,[1)))\mathfrak{e}=$
$=(1,0,[1))\mathfrak{e}\cdot(0,1,[1))\mathfrak{e}=$
$=(y,x,[0))\cdot(x,y,[0))=$
$=(y,y,[0)).$
This and the definition of the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ imply that
$(0,1,[1))\mathfrak{e}=(s,s+p,[0))\qquad\hbox{and}\qquad(1,0,[1))\mathfrak{e}=(s+p,s,[0)).$
Then for any positive integers $n_{1}$ and $n_{2}$ by usual calculations we get that
$(0,n_{1},[1))\mathfrak{e}$ $=(\underbrace{(0,1,[1))\cdot\ldots\cdot(0,1,[1))}_{n_{1}\hbox{\tiny{-times}}})\mathfrak{e}=$
$=\underbrace{(0,1,[1))\mathfrak{e}\cdot\ldots\cdot(0,1,[1))\mathfrak{e}}_{n_{1}\hbox{\tiny{-times}}}=$
$=(s,s+p,[0))^{n_{1}}=$
$=(s,s+n_{1}p,[0))$
and
$(n_{2},0,[1))\mathfrak{e}$ $=(\underbrace{(1,0,[1))\cdot\ldots\cdot(1,0,[1))}_{n_{2}\hbox{\tiny{-times}}})\mathfrak{e}=$
$=\underbrace{(1,0,[1))\mathfrak{e}\cdot\ldots\cdot(1,0,[1))\mathfrak{e}}_{n_{2}\hbox{\tiny{-times}}}=$
$=(s+p,s,[0))^{n_{2}}=$
$=(s+n_{2}p,s,[0)),$
and hence
(4) $(n_{1},n_{2},[1))\mathfrak{e}=(s+n_{1}p,s+n_{2}p,[0)).$
The definition of the natural partial order on the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ (see Proposition 4 of [5]) imply that for any positive integer $m$ we have that
$(m+1,m+1,[0))\preccurlyeq(m,m,[1))\preccurlyeq(m,m,[0)).$
Then by equalities (3), (4), and Proposition 1.4.21$(6)$ of [9] we obtain that
$(k(m+1),k(m+1),[0))$ $=(m+1,m+1,[0))\mathfrak{e}\preccurlyeq$
$\preccurlyeq(s+pm,s+pm,[0))=$
$=(m,m,[1))\mathfrak{e}\preccurlyeq$
$\preccurlyeq(m,m,[0))\mathfrak{e}=$
$=(km,km,[0)).$
The above inequalities and the definition of the natural partial order on the semigroup $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ (see Proposition 4 of [5]) imply that $km\leqslant s+pm\leqslant k(m+1)$ for any positive integer $m$. This implies that
$k\leqslant\frac{s}{m}+p\leqslant k+\frac{1}{m},$
and since $p$ is a positive integer we get that $p=k$. Hence by (4) we get that
(5) $(n_{1},n_{2},[1))\mathfrak{e}=(s+n_{1}k,s+n_{2}k,[0)),$
for all $n_{1},n_{2}\in\omega$.
It is obvious that if $s\in\{1,\ldots,k-1\}$ then $\mathfrak{e}$ is an injective monoid endomorphism of the semigroup. Hence we have that either $s=0$ or $s=k$, Simple verifications show that
$\mathfrak{e}=\left\{\begin{array}[]{ll}\gamma_{k},&\hbox{if~{}}s=0;\\ \delta_{k},&\hbox{if~{}}s=k.\end{array}\right.$
This completes the proof of the theorem. ∎
Theorem 2.
Let =$\mathscr{F}=\{[0),[1)\}$. Then for all positive integers $k_{1}$ and $k_{2}$ the following conditions hold:
  1. (1)
    $\gamma_{k_{1}}\gamma_{k_{2}}=\gamma_{k_{1}k_{2}}$;
  2. (2)
    $\gamma_{k_{1}}\delta_{k_{2}}=\gamma_{k_{1}k_{2}}$;
  3. (3)
    $\delta_{k_{1}}\gamma_{k_{2}}=\delta_{k_{1}k_{2}}$;
  4. (4)
    $\delta_{k_{1}}\delta_{k_{2}}=\delta_{k_{1}k_{2}}$.
Proof.
(1) For any $i,j\in\omega$ we have that
$(i,j,[0))\gamma_{k_{1}}\gamma_{k_{2}}$ $=(k_{1}i,k_{1}j,[0))\gamma_{k_{2}}=$
$=(k_{1}k_{2}i,k_{1}k_{2}j,[0)),$
and $(i,j,[1))\gamma_{k_{1}}=(i,j,[0))\gamma_{k_{1}}$. This implies that $\gamma_{k_{1}}\gamma_{k_{2}}=\gamma_{k_{1}k_{2}}$.
(2) Since
$(i,j,[0))\gamma_{k_{1}}\delta_{k_{2}}$ $=(k_{1}i,k_{1}j,[0))\delta_{k_{2}}=$
$=(k_{1}k_{2}i,k_{1}k_{2}j,[0)),$
and $(i,j,[1))\gamma_{k_{1}}=(i,j,[0))\gamma_{k_{1}}$ for all $i,j\in\omega$, we get that $\gamma_{k_{1}}\delta_{k_{2}}=\gamma_{k_{1}k_{2}}$.
(3) For any $i,j\in\omega$ we have that
$(i,j,[0))\delta_{k_{1}}\gamma_{k_{2}}$ $=(k_{1}i,k_{1}j,[0))\gamma_{k_{2}}=$
$=(k_{1}k_{2}i,k_{1}k_{2}j,[0)),$
and
$(i,j,[1))\delta_{k_{1}}\gamma_{k_{2}}$ $=(k_{1}(i+1),k_{1}(j+1),[0))\gamma_{k_{2}}=$
$=(k_{1}k_{2}(i+1),k_{1}k_{2}(j+1),[0)),$
and hence $\delta_{k_{1}}\gamma_{k_{2}}=\delta_{k_{1}k_{2}}$.
(4) For any $i,j\in\omega$ we have that
$(i,j,[0))\delta_{k_{1}}\delta_{k_{2}}$ $=(k_{1}i,k_{1}j,[0))\delta_{k_{2}}=$
$=(k_{1}k_{2}i,k_{1}k_{2}j,[0)),$
and
$(i,j,[1))\delta_{k_{1}}\delta_{k_{2}}$ $=(k_{1}(i+1),k_{1}(j+1),[0))\delta_{k_{2}}=$
$=(k_{1}k_{2}(i+1),k_{1}k_{2}(j+1),[0)),$
and hence $\delta_{k_{1}}\delta_{k_{2}}=\delta_{k_{1}k_{2}}$. ∎
By $\mathfrak{e}_{\boldsymbol{0}}$ we denote the annihilating monoid endomorphism of the monoid $\boldsymbol{B}_{\omega}^{\mathscr{F}}$ for the family =$\mathscr{F}=\{[0),[1)\}$, i.e., $(i,j,[p))\mathfrak{e}_{\boldsymbol{0}}=(0,0,[0))$ for all $i,j\in\omega$ and $p=0,1$. We put $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})=\boldsymbol{End}_{0}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})\setminus\{\mathfrak{e}_{\boldsymbol{0}}\}$. Theorem 2 implies that $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ is a subsemigroup of $\boldsymbol{End}_{0}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$.
Theorem 2 implies the following corollary.
Corollary 1.
If =$\mathscr{F}=\{[0),[1)\}$, then the elements $\gamma_{1}$ and $\delta_{1}$ are unique idempotents of the semigroup $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$.
Next, by ${\mathfrak{LZ}}_{2}$ we denote the left zero semigroup with two elements and by $\mathbb{N}_{u}$ the multiplicative semigroup of positive integers.
Proposition 2.
Let =$\mathscr{F}=\{[0),[1)\}$. Then the semigroup $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ is isomorphic to the direct product ${\mathfrak{LZ}}_{2}\times\mathbb{N}_{u}$.
Proof.
Put ${LZ}_{2}=\{c,d\}$. We define a map $\mathfrak{I}\colon\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})\to{\mathfrak{LZ}}_{2}\times\mathbb{N}_{u}$ by the formula
$(\mathfrak{e})\mathfrak{I}=\left\{\begin{array}[]{ll}(c,k),&\hbox{if~{}}\mathfrak{e}=\gamma_{k};\\ (d,k),&\hbox{if~{}}\mathfrak{e}=\delta_{k}.\end{array}\right.$
It is obvious that such defined map $\mathfrak{I}$ is bijective, and by Theorem 2 it is a homomorphism. ∎
Theorem 3 describes Green’s relations on the semigroup $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$. Later by $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})^{1}$ we denote the semigroup $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ with adjoined identity element.
Theorem 3.
Let =$\mathscr{F}=\{[0),[1)\}$. Then the following statements hold:
  1. (1)
    $\gamma_{k_{1}}\mathscr{R}\gamma_{k_{2}}$ in $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ if and only if $k_{1}=k_{2}$;
  2. (2)
    $\gamma_{k_{1}}\mathscr{R}\delta_{k_{2}}$ does not hold in $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ for any $\gamma_{k_{1}},\delta_{k_{2}}$;
  3. (3)
    $\delta_{k_{1}}\mathscr{R}\delta_{k_{2}}$ in $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ if and only if $k_{1}=k_{2}$;
  4. (4)
    $\gamma_{k_{1}}\mathscr{L}\gamma_{k_{2}}$ in $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ if and only if $k_{1}=k_{2}$;
  5. (5)
    $\gamma_{k_{1}}\mathscr{L}\delta_{k_{2}}$ in $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ if and only if $k_{1}=k_{2}$;
  6. (6)
    $\delta_{k_{1}}\mathscr{L}\delta_{k_{2}}$ in $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ if and only if $k_{1}=k_{2}$;
  7. (7)
    $\mathscr{H}$ is the identity relation on $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$;
  8. (8)
    $\mathfrak{e}_{1}\mathscr{D}\mathfrak{e}_{2}$ in $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ if and only if $\mathfrak{e}_{1}=\mathfrak{e}_{2}$ or there exists a positive integer $k$ such that $\mathfrak{e}_{1},\mathfrak{e}_{2}\in\{\gamma_{k},\delta_{k}\}$;
  9. (9)
    $\mathscr{D}=\mathscr{J}$ in $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$.
Proof.
(1) ()($\Rightarrow$)( ⇒ ) Suppose that $\gamma_{k_{1}}\mathscr{R}\gamma_{k_{2}}$ in $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$. Then there exist $\mathfrak{e}_{1},\mathfrak{e}_{2}\in\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})^{1}$ such that $\gamma_{k_{1}}=\gamma_{k_{2}}\mathfrak{e}_{1}$ and $\gamma_{k_{2}}=\gamma_{k_{1}}\mathfrak{e}_{2}$. The equality $\gamma_{k_{1}}=\gamma_{k_{2}}\mathfrak{e}_{1}$ and Theorem 2 imply that there exists a positive integer $p$ such that either $\mathfrak{e}_{1}=\gamma_{p}$ or $\mathfrak{e}_{1}=\delta_{p}$. In both above cases by Theorem 2 we have that
$\gamma_{k_{1}}=\gamma_{k_{2}}\mathfrak{e}_{1}=\gamma_{k_{2}}\gamma_{p}=\gamma_{k_{2}}\delta_{p}=\gamma_{k_{2}p},$
and hence $k_{2}|k_{1}$. The proof of the statement that $\gamma_{k_{2}}=\gamma_{k_{1}}\mathfrak{e}_{2}$ implies that $k_{1}|k_{2}$ is similar. Therefore we get that $k_{1}=k_{2}$.
Implication ()($\Leftarrow$)( ⇐ ) is trivial.
Statement (2) follows from Theorem 2(2).
The proof of statement (3) is similar to (1).
(4) ()($\Rightarrow$)( ⇒ ) Suppose that $\gamma_{k_{1}}\mathscr{L}\gamma_{k_{2}}$ in $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$. Then there exist $\mathfrak{e}_{1},\mathfrak{e}_{2}\in\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})^{1}$ such that $\gamma_{k_{1}}=\mathfrak{e}_{1}\gamma_{k_{2}}$ and $\gamma_{k_{2}}=\mathfrak{e}_{2}\gamma_{k_{1}}$. The equality $\gamma_{k_{1}}=\mathfrak{e}_{1}\gamma_{k_{2}}$ and Theorem 2 imply that there exists a positive integer $p$ such that $\mathfrak{e}_{1}=\gamma_{p}$. Then we have that
$\gamma_{k_{1}}=\mathfrak{e}_{1}\gamma_{k_{2}}=\gamma_{p}\gamma_{k_{2}}=\gamma_{pk_{2}},$
and hence $k_{2}|k_{1}$. The proof of the statement that $\gamma_{k_{2}}=\mathfrak{e}_{2}\gamma_{k_{1}}$ implies that $k_{1}|k_{2}$ is similar. Therefore we get that $k_{1}=k_{2}$.
Implication ()($\Leftarrow$)( ⇐ ) is trivial.
(5) ()($\Rightarrow$)( ⇒ ) Suppose that $\gamma_{k_{1}}\mathscr{L}\delta_{k_{2}}$ in $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$. Then there exist $\mathfrak{e}_{1},\mathfrak{e}_{2}\in\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})^{1}$ such that $\gamma_{k_{1}}=\mathfrak{e}_{1}\delta_{k_{2}}$ and $\delta_{k_{2}}=\mathfrak{e}_{2}\gamma_{k_{1}}$. The equality $\gamma_{k_{1}}=\mathfrak{e}_{1}\delta_{k_{2}}$ and Theorem 2 imply that there exists a positive integer $p$ such that $\mathfrak{e}_{1}=\gamma_{p}$. Then we have that
$\gamma_{k_{1}}=\mathfrak{e}_{1}\delta_{k_{2}}=\gamma_{p}\delta_{k_{2}}=\gamma_{pk_{2}},$
and hence $k_{2}|k_{1}$. The equality $\delta_{k_{2}}=\mathfrak{e}_{2}\gamma_{k_{1}}$ and Theorem 2 imply that there exists a positive integer $q$ such that $\mathfrak{e}_{1}=\delta_{q}$. Then we have that
$\delta_{k_{2}}=\mathfrak{e}_{2}\gamma_{k_{1}}=\delta_{q}\gamma_{k_{1}}=\gamma_{qk_{1}},$
and hence $k_{1}|k_{2}$. Thus we get that $k_{1}=k_{2}$.
Implication ()($\Leftarrow$)( ⇐ ) is trivial.
(6) ()($\Rightarrow$)( ⇒ ) Suppose that $\delta_{k_{1}}\mathscr{L}\delta_{k_{2}}$ in $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$. Then there exist $\mathfrak{e}_{1},\mathfrak{e}_{2}\in\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})^{1}$ such that $\delta_{k_{1}}=\mathfrak{e}_{1}\delta_{k_{2}}$ and $\delta_{k_{2}}=\mathfrak{e}_{2}\delta_{k_{1}}$. The equality $\delta_{k_{1}}=\mathfrak{e}_{1}\delta_{k_{2}}$ and Theorem 2 imply that there exists a positive integer $p$ such that $\mathfrak{e}_{1}=\delta_{p}$. Then we have that
$\delta_{k_{1}}=\mathfrak{e}_{1}\delta_{k_{2}}=\delta_{p}\delta_{k_{2}}=\delta_{pk_{2}},$
and hence $k_{2}|k_{1}$. The proof of the statement that $\delta_{k_{2}}=\mathfrak{e}_{2}\delta_{k_{1}}$ implies that $k_{1}|k_{2}$ is similar. Hence we get that $k_{1}=k_{2}$.
Implication ()($\Leftarrow$)( ⇐ ) is trivial.
(7) By statements (1), (2), and (3), $\mathscr{R}$ is the identity relation on the semigroup $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$. Then so is $\mathscr{H}$, because $\mathscr{H}\subseteq\mathscr{R}$.
Statement (8) follows from statements (1)–(6).
(9) Suppose to the contrary that $\mathscr{D}\neq\mathscr{J}$ in $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$. Since $\mathscr{D}\subseteq\mathscr{J}$, statement (8) implies that there exist $\mathfrak{e}_{1},\mathfrak{e}_{2}\in\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})^{1}$ such that $\mathfrak{e}_{1}\mathscr{J}\mathfrak{e}_{2}$ and $\mathfrak{e}_{1},\mathfrak{e}_{2}\notin\{\gamma_{k},\delta_{k}\}$ for any positive integer $k$. Then there exist distinct positive integers $k_{1}$ and $k_{2}$ such that $\mathfrak{e}_{1}\in\{\gamma_{k_{1}},\delta_{k_{1}}\}$ and $\mathfrak{e}_{2}\in\{\gamma_{k_{2}},\delta_{k_{2}}\}$. Without loss of generality we may assume that $k_{1}<k_{2}$. Since $\mathfrak{e}_{1}\mathscr{J}\mathfrak{e}_{2}$ there exist $\mathfrak{e}_{1}^{\prime},\mathfrak{e}_{2}^{\prime},\mathfrak{e}_{1}^{\prime\prime},\mathfrak{e}_{2}^{\prime\prime}\in\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})^{1}$ such that $\mathfrak{e}_{1}=\mathfrak{e}_{1}^{\prime}\mathfrak{e}_{2}\mathfrak{e}_{1}^{\prime\prime}$ and $\mathfrak{e}_{2}=\mathfrak{e}_{2}^{\prime}\mathfrak{e}_{2}\mathfrak{e}_{2}^{\prime\prime}$. Since $\mathfrak{e}_{1}\in\{\gamma_{k_{1}},\delta_{k_{1}}\}$ and $\mathfrak{e}_{2}\in\{\gamma_{k_{2}},\delta_{k_{2}}\}$, the equality $\mathfrak{e}_{1}=\mathfrak{e}_{1}^{\prime}\mathfrak{e}_{2}\mathfrak{e}_{1}^{\prime\prime}$, Theorems 1 and  2 imply that $k_{2}|k_{1}$. This contradicts the inequality $k_{1}<k_{2}$. The obtained contradiction implies the requested statement ∎
Remark 4.
Since $\mathfrak{e}_{\boldsymbol{0}}$ is zero of the semigroup $\boldsymbol{End}_{0}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$, the classes of equivalence of Green’s relations of non-zero elements of $\boldsymbol{End}_{0}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ in the semigroup $\boldsymbol{End}_{0}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$ coincide with their corresponding classes of equivalence in $\boldsymbol{End}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$, and moreover we have that
$\boldsymbol{L}_{\mathfrak{e}_{\boldsymbol{0}}}=\boldsymbol{R}_{\mathfrak{e}_{\boldsymbol{0}}}=\boldsymbol{H}_{\mathfrak{e}_{\boldsymbol{0}}}=\boldsymbol{D}_{\mathfrak{e}_{\boldsymbol{0}}}=\boldsymbol{J}_{\mathfrak{e}_{\boldsymbol{0}}}=\{\mathfrak{e}_{\boldsymbol{0}}\}$
in the semigroup $\boldsymbol{End}_{0}^{*}(\boldsymbol{B}_{\omega}^{\mathscr{F}})$.
Acknowledgements
The authors acknowledge the referee for his/her comments and suggestions.